OpenStudy (anonymous):

When a trinomial is factored as (x + m)(x + n), what is the product of m and n? A.The coefficient of the x-term of the trinomial B.The constant term of the trinomial C.The degree of the trinomial D.The coefficient of the x2-term of the trinomial

OpenStudy (mathstudent55):

Use FOIL: (x + m)(x + n) = x^2 + nx + mx + mn Compare with ax^2 + bx + c: x^2 = x^2 nx + mx = bx mn = c, the constant term of the trinomial

OpenStudy (anonymous):

so it would be B @mathstudent55

OpenStudy (mathstudent55):

correct